An inductor is hooked up to an AC voltage source. The voltage source has EMF V0 and frequency f. The current amplitude in the inductor is I0. What is the reactance XL of the inductor? What is the inductance L of the inductor?

An rms voltage of 19.5 V with a frequency of 1.20 kHz is applied to a 0.365 μF capacitor. What is the rms current in this circuit? By what factor does the current change if the frequency of the voltage is doubled?Calculate the current for a frequency of 2.40 kHz .
Answer #1

XL = Vo/Io

for inductance

2 pi f L = Vo/Io

L = Vo / ( 2 pi f Io)

======

i = 19.5 /( 1 / ( 2*3.14* 1200* 0.365* 10^-6))

i = 0.0536 A

current increases on increasing frequency,

so,

i' = 2 i = 0.1073 A

=======

Comment before rate in case any doubt, will reply for sure.. goodluck